Solving a System of Linear Equations by Elimination/Subtraction
Can you help me solve this equation? Also, can you explain it too?
3x-3y=-6
-5x+6y=12
Thank you!!!

Answer by mananth(16949)

3x-3y=-6....................1
-5x+6y=12....................2
...
if you multiply the first equation by 2 then you will get
6x-6y=-12
when you add up this equation to equation 2 "y" will get eliminated
we get
-5x+6y+6x-6y=12-12
x=0
...
plug value of x in equation 1
3*0-3y=-6
-3y=-6
/-3
y=-6/-3
y=2
...
m.ananth@hotmail.ca